OpenStudy (anonymous):

How can x + 8 = -x + 7 be set up as a system of equations?

OpenStudy (mrhoola):

solve for X first. x= -1/2 then plug in X to one side but generate two equations x+8 = -1/2 +7 (eq1) -x+7 =-1/2 +8 (eq2) That is your system of two linear equations . you can verify it by solving the system

OpenStudy (arabpride):

Correct! :3 Please close this post if you're done. Thanks.